OpenStudy (anonymous):

factor x^2+x-12 explain how to do please

OpenStudy (earthcitizen):

http://www.wolframalpha.com/input/?i=+x%5E2%2Bx-12

jhonyy9 (jhonyy9):

x^2 +4x -3x -4*3

jhonyy9 (jhonyy9):

(x-3)(x+4)

OpenStudy (earthcitizen):

to explain you might need to use quadratic formula

OpenStudy (anonymous):

There are a few characteristics of the equation you entered to notice. First, you only have one variable. Next, the highest exponent in your equation is a 2. That means you have a quadratic equation. Most of the quadratic equations that you incounter in school are purposely written so that they are the result of FOIL (first, outer, inner, last). In your example that means that we multiply out (x-3)(x+4) like this. first=(x)(x) outer=(-3)(4) inner = (-3)(x) and last = (4)(x). Now you add up all those terms. x^2 + 4x - 3x - 12. Combine like terms and you'll have your equation. x^2 + x - 12. So when they asked you to factor they want you to find terms that will multiply together to give you the equation. Since you don't have a coefficient on the x^2 term you can find it by listing the factors of -12. (1,2,3,4,6,12) but notice that you will have to have one of them be negative to get -12. Now look at those numbers and try to find to that will multiply to equal -12, but add to equal the coefficient of your x term in this case (1). That will be 4 and -3. Another way to find the numbers you need to work with is to run them through the quadratic formula. It will be easy to find in a web search. "a" is the coefficient of your x^2 term. "b" is the coefficient of you x term. "c" is the content in your equation. If you don't have an x term its coefficient is 0.